OpenStudy (anonymous):

Which choice is the explicit formula for the following geometric sequence? 0.5, -0.1, 0.02, -0.004, 0.0008, ... A. an = 0.5(0.2)^n B. an = 0.5(-0.2)^n-1 C. an = -0.5(0.3)^n-1 D. an = -0.5(0.2)^n-1 E. an = -0.5(-0.2)^n-1

OpenStudy (anonymous):

@Godlovesme could you help?

OpenStudy (anonymous):

im lost on how to find the answer on this..

OpenStudy (godlovesme):

me too lol

OpenStudy (anonymous):

lol do you know anyone who could help us? cause this stuff is blah. xD

OpenStudy (godlovesme):

fr!!! C: @sammixboo @Loser66 help?

OpenStudy (anonymous):

well the method to get to the answer. not neccisarly give it to me. but help me get to it and guide me on what i do wrong

OpenStudy (loser66):

0.5 is the \(1^{st}\) number, right? that is n =1

OpenStudy (loser66):

so, \(a_1 = 0.5\)

OpenStudy (anonymous):

yes

OpenStudy (loser66):

now, I test the first option: \(a_1= 0.5 * (0.2)^1 =??\) pretty sure that for the first option, \(a_1\) can't be 0.5, right?

OpenStudy (anonymous):

Ok yea?

OpenStudy (loser66):

Now, I test the second option: \(a_1 = 0.5*(-0.2)^{1-1}=0.5*(-0.2)^0=0.5\) oh yeah... sounds good

OpenStudy (loser66):

No need to test the rest, since for n =1, their \(a_1) = - 0.5, so, they are not the correct answer.

OpenStudy (anonymous):

sorry my internet went to poop..

OpenStudy (anonymous):

so the 0.5 and 0.2 will both be negative? I am a little confused...

OpenStudy (loser66):

All you need is doing the same for the second term \(a_2\) with the second option to make sure that you are correct when choosing it.

OpenStudy (loser66):

the given second term is -0.1 now , test the formula of b) it says \(a_2 = 0.5*(-0.2)^{2-1}=0.5*(-0.2)^1=0.5*(-0.2)=-0.1\) bingo